OpenStudy (anonymous):

Why is the integral of dx 1 but the intgral of dx ^ n not possible

myininaya (myininaya):

what

OpenStudy (anonymous):

\[\int\limits_{}^{}dx = 1 \] but \[\int\limits_{}^{}dx ^{3} \] does not = 1^3

OpenStudy (anonymous):

Why can dx not be integrated if it has an exponent higher than 1?

OpenStudy (anonymous):

dx is not a term, or a variable, or anything other than an identifier that we are trying to integrate something. There is no function "dx"; when you see something like \[\int\limits_{}^{} dx\] it really means \[\int\limits_{}^{} 1 dx\] Thus you're integrating 1.

OpenStudy (anonymous):

right, so If I come to a dx^3 by u-substitution shouldnt it just be like integrating 1^3

OpenStudy (anonymous):

No, because dx itself is not a function that you can raise to a power. If you integrate \[\int\limits_{}^{} 1^{3} dx\] you are still integrating 1. You are not integrating "dx"

OpenStudy (anonymous):

You shouldn't come to a \[dx^3\]by u-substitution in the first place.